# Practitioners’ Challenges and Perceptions of CI Build Failure Predictions at Atlassian

### v1

@@ 5.2. Implication for Practitioners
Optimizing CI Processes with CI Build Prediction: The results of our survey indicate that developers appreciate the CI build prediction as a proactive measure for catching issues in builds. It’s recommended that organizations integrate such technique into their CI processes. This technique not only identifies potential issues before code merging but also aids in decision-making, especially in fast-paced development environments where the cost of a delayed reaction can be considerable.
Balancing Reliance with Human Judgment: While CI build prediction are valuable, there’s a risk of developers becoming overly reliant on these techniques. It’s recommended that while these techniques should be used to aid the review process, they should not replace the need for thorough human review. Practitioners should foster an environment where techniques are used as a complement to human expertise.
Encouraging Learning from CI Build Failures: Although CI build failures can be disruptive, they are also seen as opportunities for learning and improving build processes. Practitioners should draw lessons from these failures. This involves not just fixing the immediate issues but also understanding the underlying causes and taking steps to prevent similar issues in the future.
Improvement of Code Review Practices: Aligning with the findings of prior research (Rahman and Roy, 2017; Zampetti et al., 2019; Cassee et al., 2020), our study confirms that meaningful and constructive code review practices are essential for a successful CI process. Practitioners should encourage teams to engage in thorough code reviews, focusing on the substance of code changes. Investing time in comprehensive reviews can preempt issues that lead to build failures.
@@ 6. Threats To Validity
Internal validity concerns the quality of the study design and the rigorousness of its execution. In our study, these threats are related to data mining, model construction, as well as survey design and analysis. We directly used the data from the Atlassian database while complying with their privacy policies, providing us with access to the latest and most relevant dataset. To enhance data accuracy, we conducted a manual inspection and filtered out the apparent outliers from the dataset. The selection of factors for the LRM model may be a potential threat to our study. We mitigated this by relying on the metrics previously established and validated by the research community. Other factors (e.g., human factors) may also play a role in CI build outcomes. However, the selection of each factor is governed by the limitation of data availability. Thus, we cannot analyze such factors due to limited access. In designing our survey, we created questions to be both clear and easy to understand. While acknowledging some potential biases in analyzing open-ended questions, we addressed this by manually conducting a thematic analysis approach for unbiased analysis and engaging multiple coders to ensure an accurate interpretation of the data. In addition, the developers facing significant challenges in CI builds may be more motivated to participate in surveys. To address this bias, we invite a wide range of developers within Atlassian, not solely those who experienced CI build failures.
@@ 7. Related Work
In this section, we discuss the related work.
7.1. CI Build Failures
@@
CI build failures can disrupt the development work of the entire team for a considerable time (Kerzazi et al., 2014). Such long build times greatly increase the cost of human and computing resources, and hence become a common barrier faced by software organizations adopting CI (Hilton et al., 2017). To address CI build failures, many studies aim to understand the underlying causes. Seo et al. (Seo et al., 2014) explored the compiler errors that occur in the build process. They developed a classifier of compilation errors leading to build failures in Java and C++ environments. Kerzazi et al. (Kerzazi et al., 2014) examined the impact of socio-technical congruence on build failures. They observed a high percentage of build failures (17.9%) that brings a potential cost of about 2,035 man-hours, considering that each failure needs one hour of work to succeed. Rausch et al. (Rausch et al., 2017) and Beller et al. (Beller et al., 2017a) focused on the impact of failed test cases on CI build. They highlight the importance of testing in CI build failure. Orellana Cordero et al. (Orellana et al., 2017) extended this focus to test-related build failures in open-source projects. Different from prior studies, our work focuses on investigating CI build failures within the context of Atlassian’s development environment. We explore both technical factors and the human and organizational factors that potentially influence CI build failures. This enables us to uncover a broader spectrum of potential causes and impacts of CI build failures.
7.2. CI Build Prediction
@@
Recent works leverage various techniques to predict the outcomes of CI builds. Xia and Li (Xia and Li, 2017) explored 6 predictive models across 126 open-source projects to predict the CI build outcome. Finlay et al. (Finlay et al., 2014) used data stream mining techniques based on code metrics (i.e., basic metrics, dependency metrics, complexity metrics, cohesion metrics, and Halstead metrics) to predict build outcomes. Ni and Li (Ni and Li, 2017) predicted build outcomes in CI based on file-level metrics from the current and previous builds and failure statistics from historical builds. Hassan and Wang (Hassan and Wang, 2017) leveraged metrics from the current and previous build. Chen et al. (Chen et al., 2020) also extract the metrics from historical builds for their prediction. Our findings further emphasize the vital role that factors related to historical builds have in predicting CI build failures. This confirms that these prior findings derived from open-source projects are applicable to the industrial context of Atlassian. In addition, some researchers (Wolf et al., 2009; Schröter, 2010) integrated social network analysis to capture communication structure. They incorporate these structures into models to predict build outcomes.
@@
Despite recent advances, build outcome prediction still faces challenges that limit the practical application of these techniques in CI. Firstly, the explanation of the predictive model needs to be investigated. The effectiveness of the predictive model hinges on their ability to provide insightful explanations (Tantithamthavorn and Jiarpakdee, 2021; Jiarpakdee et al., 2020, 2021). To help developers effectively address CI build failures, it is necessary to explain the models of these CI build predictions and understand the reasoning behind the predictions. Secondly, there’s a gap in understanding how developers perceive and interact with the CI build prediction techniques. Grasping developers’ expectations and concerns is crucial for the successful integration and adoption of these techniques in real-world scenarios. Thus, we recognize our work as a complement to those previous works.

### v2

@@ 5.2. Implications for Practitioners
Encouraging Learning from CI Build Failures: RQ2 shows that although CI build failures can be disruptive, they are also seen as opportunities for learning from failures and improving build processes. Practitioners should draw lessons from these failures. This involves not just fixing the immediate issues but also understanding the underlying causes and taking steps to prevent similar issues in the future.
Improvement of Code Review Practices: Aligning with the findings of prior research (Rahman and Roy, 2017; Zampetti et al., 2019; Cassee et al., 2020), our study confirms that meaningful and constructive code review practices are essential for a successful CI process based on the findings of RQ2. Practitioners should encourage teams to engage in thorough code reviews, focusing on the substance of code changes. Investing time in comprehensive reviews can preempt issues that lead to build failures.
Optimizing CI Processes with CI Build Prediction: The results of RQ3 indicate that developers appreciate the CI build prediction as a proactive measure for catching issues in builds. It’s recommended that organizations integrate such technique into their CI processes. This technique not only identifies potential issues before code merging but also aids in decision-making, especially in fast-paced development environments where the cost of a delayed reaction can be considerable.
Balancing Reliance with Human Judgment: From the results of RQ3, we found that while CI build prediction are valuable, there’s a risk of developers becoming overly reliant on these techniques. It’s recommended that while these techniques should be used to aid the review process, they should not replace the need for thorough human review. Practitioners should foster an environment where techniques are used as a complement to human expertise.
@@ 6. Threats To Validity
Internal validity concerns the quality of the study design and the rigorousness of its execution. In our study, these threats are related to data mining, model construction, as well as survey design and analysis. We directly utilized the aggregated data sourced from Atlassian’s developer database (excluding any of Atlassian’s customers’ data). This provided us with access to the latest and most relevant dataset. To enhance data accuracy, we conducted a manual inspection and filtered out the apparent outliers from the dataset. The selection of factors for the LRM model may be a potential threat to our study. We mitigated this by relying on the metrics previously established and validated by the research community. Other factors (e.g., human factors) may also play a role in CI build outcomes. However, the selection of each factor is governed by the limitation of data availability. Thus, we cannot analyze such factors due to limited access. In designing our survey, we created questions to be both clear and easy to understand. While acknowledging some potential biases in analyzing open-ended questions, we addressed this by manually conducting a thematic analysis approach for unbiased analysis and engaging multiple coders to ensure an accurate interpretation of the data. In addition, the developers facing significant challenges in CI builds may be more motivated to participate in surveys. To address this bias, we invite a wide range of developers within Atlassian, not solely those who experienced CI build failures.
@@ 7. Related Work
In this section, we discuss the related work related to CI build failures and CI build predictions.
@@
CI Build Failures. CI build failures can disrupt the development work of entire teams for a considerable time (Kerzazi et al., 2014; Maipradit et al., 2023). Such long build times greatly increase the cost of human and computing resources, and hence become a common barrier faced by software organizations adopting CI (Hilton et al., 2017). To address CI build failures, many studies aim to understand the underlying causes (Fu et al., 2024). Seo et al. (Seo et al., 2014) explored compiler errors that occur in the build process. They developed a classifier of compilation errors leading to build failures. Kerzazi et al. (Kerzazi et al., 2014) examined the impact of socio-technical congruence on build failures. They observed a high percentage of build failures (17.9%) that brings a potential cost of about 2,035 man-hours, considering that each failure needs one hour of work to succeed. Jin et al. (Jin and Servant, 2022) developed a set of rules that indicates the safe builds to skip. Similarly, Abdalkareem et al. (Abdalkareem et al., 2019) manual investigate 588 java projects. They proposed the rules to identify the CI skipped commit. Rausch et al. (Rausch et al., 2017) and Beller et al. (Beller et al., 2017a) focused on the impact of failed test cases on CI build. They highlight the importance of testing in CI build failure. Orellana Cordero et al. (Orellana et al., 2017) extended this focus to test-related build failures in open-source projects. Gallaba et al. (Gallaba et al., 2018) investigate 3.7 million build jobs to analyze the noise and heterogeneity in build outcome data.
Different from prior studies, our work focuses on investigating CI build failures within the context of Atlassian’s development environment. We explore both technical factors and the human and organizational factors that potentially influence CI build failures. This enables us to uncover a broader spectrum of potential causes and impacts of CI build failures.
@@
CI Build Predictions. Recent works leverage various techniques to predict the outcomes of CI builds. Xia and Li (Xia and Li, 2017) explored 6 predictive models across 126 open-source projects to predict the CI build outcome. Finlay et al. (Finlay et al., 2014) used data stream mining techniques based on code metrics (i.e., basic metrics, dependency metrics, complexity metrics, cohesion metrics, and Halstead metrics) to predict build outcomes. Jin et al. (Jin and Servant, 2020) proposed SmartBuildSkip to predict the first build in a sequence of build failures. Abdalkareem et al. (Abdalkareem et al., 2020) proposed a machine learning approach to determine the commits that can be CI skipped. Barrak et al. (Barrak et al., 2021) performed a replication study using different datasets and features to explain and predict the outcome of builds. Jin and Servant (Jin and Servant, 2023) introduce a hybrid technique called HybridCISave to predict CI build failure. Ni and Li (Ni and Li, 2017) predicted build outcomes in CI based on file-level metrics from the current and previous builds and failure statistics from historical builds. Hassan and Wang (Hassan and Wang, 2017) leveraged metrics from the current and previous build. Chen et al. (Chen et al., 2020) also extract the metrics from historical builds for their prediction. Our findings further emphasize the vital role that factors related to historical builds have in predicting CI build failures. This confirms that these prior findings derived from open-source projects are applicable to the industrial context of Atlassian. In addition, some researchers (Wolf et al., 2009; Schröter, 2010) integrated social network analysis to capture communication structure. They incorporate these structures into models to predict build outcomes.
@@
Despite recent advances, build outcome prediction still faces challenges that limit the practical application of these techniques in CI. Firstly, the explanation of the predictive model needs to be investigated. The effectiveness of the predictive model hinges on their ability to provide insightful explanations (Tantithamthavorn and Jiarpakdee, 2021; Jiarpakdee et al., 2020, 2021; Tantithamthavorn et al., 2015; Pornprasit et al., 2021; Tantithamthavorn et al., 2020, 2021, 2023). To help developers effectively address CI build failures, it is necessary to explain the models of these CI build predictions and understand the reasoning behind the predictions. Secondly, there’s a gap in understanding how developers perceive and interact with the CI build prediction techniques. Grasping developers’ expectations and concerns is crucial for the successful integration and adoption of these techniques in real-world scenarios. Thus, we recognize our work as a complement to those previous works.
@@ 8. Conclusion
9. Disclaimer
The perspectives and conclusions presented in this document are solely the authors’ and should not be interpreted as representing the official policies or endorsements of Atlassian or any of its subsidiaries and affiliates. Additionally, the outcomes of this study are independent of, and should not be construed as an assessment of, the quality of products offered by Atlassian.